OpenStudy (anonymous):

Solve |2x + 8| = 20.

OpenStudy (anonymous):

{-14, 0} {-14, 6} {-6, 14} {0, 14} but these are the answers what am i doing wrong?

OpenStudy (raden):

20-8=12

OpenStudy (anonymous):

im confused

OpenStudy (jziggy):

You need to subtract 8 from both sides

OpenStudy (anonymous):

ok i got 2x=12

OpenStudy (raden):

so, x=12/2 = ... ?

OpenStudy (anonymous):

6

OpenStudy (raden):

correct, fot the first case!

OpenStudy (raden):

*for

OpenStudy (anonymous):

then i plug in 6 for x?

OpenStudy (jziggy):

Since the equation has an absolute value in it you need to set up 2 different equations. One for each sign.

OpenStudy (anonymous):

what chu mean ziggy?

OpenStudy (jziggy):

(2x+8) = 20 and -(2x + 8) = 20

OpenStudy (anonymous):

???

OpenStudy (raden):

yeah, that's i mean u just solved the first case : 2x+8=20, u got x=6 but not yet solve for -(2x+8)=20, find this solution!

OpenStudy (jziggy):

I'm not sure how you factored out your x, the second step for that should be to add 8 to both sides of the equation.

OpenStudy (raden):

dont use x=6 to this problem (it is was clear) just solve for x

OpenStudy (jziggy):

Yeah, there are two solutions for x. 6 is the first but you need to find the second of them.

OpenStudy (anonymous):

i got -6 now?

OpenStudy (raden):

still incorrect, try again

OpenStudy (jziggy):

-(2x + 8) = 20 -2x - 8 = 20 + 8 + 8 -2x = 28

OpenStudy (anonymous):

im over this im going to guess thanks tho!!

OpenStudy (raden):

so, what's ur answer Jessica ?

OpenStudy (anonymous):

D

OpenStudy (jziggy):

You know that one of the answers is 6, that narrows it down to one choice anyways.

OpenStudy (anonymous):

B

OpenStudy (jziggy):

Correct

OpenStudy (anonymous):

can you help me out with one more like this?

OpenStudy (anonymous):

ill appreciate it =)

OpenStudy (jziggy):

Sure

OpenStudy (anonymous):

Solve |x + 9| < 12..

OpenStudy (anonymous):

got it

OpenStudy (jziggy):

Oh, ok *erases the already typed work*